# Video Understanding with Large Language Models: A Survey

### v7

@@ I-A1 Conventional Methods
In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) [1], Speeded-Up Robust Features (SURF) [2], and Histogram of Oriented Gradients (HOG) [3] were used to capture key information in videos. Background Subtraction [4], optical flow methods [5], and Improved Dense Trajectories (IDT) [6, 7] were used to model the motion information for tracking. Since videos can be viewed as time series data, temporal analysis techniques such as Hidden Markov Models (HMM) [8] have also been used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) [9], Decision Trees [10], and Random Forests were also used in video classification and recognition tasks. Cluster analysis [11] for classifying video segments, or Principal Component Analysis (PCA) [12, 13] for data dimensionality reduction have also been commonly used methods for video analysis.
@@ I-A2 Early Neural Video Models
Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities. DeepVideo [14] and [15] were early methods introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted feature method due to the inadequate use of motion information. Two-stream networks [16] combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted [17]. Temporal Segment Network (TSN) [18] was also designed for long-form video understanding by analyzing and aggregating video segments. Besides TSN, Fisher Vectors (FV) encoding [19], Bi-Linear encoding [20], and Vector of Locally Aggregated Descriptors (VLAD) [21] encoding were introduced [22]. These methods improved performance on the UCF-101 [23] and HMDB51 [24] datasets. Unlike two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) [25]. Inflated 3D ConvNets (I3D) [26] utilizes the initialization and the architecture of 2D CNN, Inception [27], to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) [28] and Something-Something [29] datasets to evaluate the model’s performance in more challenging scenarios. ResNet [30], ResNeXt [31], and SENet [32] were also adapted from 2D to 3D, resulting in the emergence of R3D [33], MFNet [34], and STC [35]. To improve the efficiency, the 3D convolution has been decomposed into cascade 2D and 1D convolution in various studies (e.g., S3D [36], ECO [37], P3D [38]). LTC [39], T3D [40], Non-local [41], and V4D [42] focus on long-form temporal modeling, while CSN [43], SlowFast [44], and X3D [45] tend to attain high efficiency. The introduction of Vision Transformers (ViT) [46] promotes a series of prominent models (e.g., TimeSformer [47], VidTr [48], ViViT [49], MViT [50]).
@@ I-A3 Self-supervised Video Pretraining
Transferability [51, 52] in self-supervised pretraining models [53] for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ requirements for extensive task-specific data. VideoBERT [54] is an early attempt to perform video pretraining. Based on the bidirectional language model BERT [55], pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pretrained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the “pretraining-finetuning” paradigm, many studies on pretrained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT [56], SpatiotemporalMAE [57], OmniMAE [58], VideoMAE [59], MotionMAE [60]) or training strategies (MaskFeat [61], VLM [62], ALPRO [63], All-in-One transformer [64], MaskViT [65], CLIP-ViP [66], Singularity [67], LF-VILA [68], EMCL [69], HiTeA [70], CHAMPAGNE [71]).
@@ I-A4 Large Language Models for Video Understanding
Recently, large language models (LLMs) have advanced rapidly [72]. The emergence of large language models pretrained on extensive datasets has introduced a novel in-context learning capability [73]. This allows them to handle various tasks using prompts without the need for fine-tuning. ChatGPT [74] is the first groundbreaking application built on this foundation. This includes capabilities like generating code and invoking tools or APIs of other models for their use. Many studies are exploring using LLMs like ChatGPT to call vision models APIs to solve the problems in the computer vision field, including Visual-ChatGPT [75]. The advent of instruct-tuning has further enhanced these models’ ability to respond effectively to user requests and perform specific tasks [76]. LLMs integrated with video understanding capabilities offer the advantage of more sophisticated multimodal understanding, enabling them to process and interpret complex interactions between visual and textual data. Similar to their impact in Natural Language Processing (NLP) [77], these models act as more general-purpose task solvers, adept at handling a broader range of tasks by leveraging their extensive knowledge base and contextual understanding acquired from vast amounts of multimodal data. This allows them to not only understand visual content but also reason about it in a way that is more aligned with human-like understanding. Many works also explore using LLMs in video understanding tasks, namely, Vid-LLMs.
@@ I-B Related Surveys
Previous survey papers either study specific sub-tasks in the area of video understanding or focus on methodologies beyond video understanding. For example, [78] surveys multimodal foundation models for general vision-language tasks, which includes both image and video applications. [79] and [80] focus on surveying video captioning and action recognition tasks, respectively. Other video understanding tasks, such as the video question answering and grounding, are not considered. Moreover, [81], [82], and [77] survey video-related methodologies, such as video diffusion models and LLMs, lacking the concentration on video understanding. [83] centers primarily on video foundation models, with insufficient attention given to language model-based approaches. Despite the significant value to the community, previous survey papers leave a gap in surveying the general video understanding task based on large language models. This paper fills this gap by comprehensively surveying the video understanding task using large language models.
@@ II-A1 Abstract Understanding Tasks
  • Video Classification & Action Recognition: Video classification and action recognition classify videos based on class labels or activities and events categories within a video sequence. Datasets specifically introduced for these tasks include UCF-101 [23], HMDB51 [24], Hollywood [84], ActivityNet [85], Charades [86], Kinetics-400 [28], Kinetics-600 [87], Kinetics-700 [88], SomethingSomethingV2 [29], HACS [89], YouTube8M [90], and PortraitMode-400 [91]. Usually, Top-K accuracy is adopted as the main metric for these tasks.
  • Text-Video Retrieval: Text-video retrieval task matches and retrieves relevant video clips based on the similarity between video clips and the input textual descriptions. Datasets like Kinetic-GEB [92], MSRVTT [93], DiDeMo [94], YouCook2 [95], and Youku-mPLUG [96] are relevant to this task. The standard evaluation metric for this task is Recall at K (R@K), which measures the accuracy of the first K retrieved results.
  • Video-to-Text Summarization: Video-to-text summarization is a task that generates concise textual summaries of videos. The video summarization approaches are trained to extract and interpret key visual and audio content to produce coherent and informative summaries. ViTT [97], VideoXum [98], VideoInstruct-100K [99], and Instrcut-V2Xum [100] are datasets that related to the task. Metrics of BLEU, METEOR, CIDEr, and ROUGE-L often evaluate this task.
  • Video Captioning: Video captioning generates descriptive and coherent textual captions of given videos. The video caption models usually use visual and auditory information from video to produce accurate and contextually relevant descriptions. Notable datasets for this tasks are MSVD [101], MSR-VTT [102], TGIF [103], Charades [86], Charades-Ego [104], YouCook2 [95], Youku-mPLUG [96], VAST-27M [105], and VideoInstruct-100K [99]. This task is often evaluated by metrics of BLEU, METEOR, CIDEr, and ROUGE-L.
  • Video QA: Video Question-Answering (VQA) aims to answer textual questions based on a given video, where the model analyzes visual and auditory information, understands the context, and eventually generates accurate responses. Datasets involved in the QA task are VCR [106], MSVD-QA [93], MSRVTT-QA [93], TGIF-QA [107], Pororo-QA [108], TVQA [109], ActivityNet-QA [110], and NExT-QA [111]. This task is evaluated using Top-1, Top-K accuracy.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content. F1-score, Spearman, and Kendall usually evaluate this task as metrics. Commonly-used datasets include SumMe [112], TVSum [113], Ads-1k [114], VideoXum [98], Instrcut-V2Xum [100].
@@
  • Video Highlight Detection: Video highlight detection aims at identifying and extracting the most important and interesting segments from a video. Commonly used datasets on this task include the YouTube Highlights [115], the TV-Sum [113], and the Videos Titles in the Wild (VTW) [116].
@@
  • Temporal Action/Event Localization: This task aims at identifying the precise temporal segments of actions or events within a video. By analyzing sequential frames, models trained for this task must indicate when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [117], ActivityNet-1.3 [85], and UnAV-100 [118].
@@
  • Temporal Action Proposal Generation: Temporal action proposal generation involves generating candidate segments within a video that are likely to contain actions or events. Relevant datasets like THUMOS’14 [117], ActivityNet [85], and Charades [86] are used for the training and evaluation for this task.
@@
  • Video Temporal Grounding: Video temporal grounding is the task of locating specific moments or intervals within a video that correspond to a given textual query. This process involves aligning linguistic descriptions with visual content, enabling precise identification of relevant segments for applications in video search and content analysis. Common benchmarks are Charades-STA [119], ViTT [97], DiDeMo [94], and PU-VALOR [120]. The metrics of R1@0.5 and R1@0.7 often evaluate this task.
@@
  • Moment Retrieval: Moment retrieval is the task of identifying and extracting precise video segments that correspond to a given textual or visual query, which aligns semantic content between queries and video frames. DiDeMo [94] is a dataset for this task.
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying certain frames in a video where significant changes occur and splitting videos based on different events or activities. Kinetics-GEBD [121] is a widely-used dataset for this task.
@@
  • Generic Event Boundary Captioning & Grounding: Generic event boundary captioning and grounding involve identifying and describing the transition points between significant events in a video, where Kinetics-GEB+ [92] is the dataset for this task.
@@
  • Dense Video Captioning: Dense video captioning [122, 123, 124, 125, 126] aims at generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [127], VidChapters-7M [128], YouCook2 [95], and ViTT [97].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking aims at continuously identifying and following the position of specific objects within a video over time. A good tracking model should maintain accurate and consistent trajectories of objects, even for videos with occlusions, changes in appearance, and motions. Benchmarks like OTB [129], UAV [130], and VOT [131] are commonly used for this task.
  • Re-Identification: Re-Identification (ReID) is the task of recognizing and matching individuals or objects across different video frames or camera views. Common datasets in ReID are Market-1501 [132], CUHK03 [133], MSMT17 [134], and DukeMTMC-reID [135].
  • Video Saliency Detection: Video saliency detection aims at identifying the most visually important and attention-grabbing regions in a video [136]. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [137], Hollywood-[86], UCF-Sports [138], AVAD [139], Coutrot1 [140], Coutrot2 [141], ETMD [142], and SumMe [112].
  • Video Object Segmentation: Video object segmentation aims at partitioning a video into segments that correspond to individual objects, accurately delineating their boundaries over time. YouTube-VOS [143] and DAVIS [144] are datasets related to this task.
  • Video Instance Segmentation: Video instance segmentation is the task of identifying, segmenting, and tracking each unique object instance within a video. YouTube-VIS [145] and Cityscapes-Seq [146] are two common benchmarks for this task.
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [147] is a common benchmark for this task.
  • Spatiotemporal Grounding: Spatiotemporal grounding aims to identify and localize specific objects or events within a video’s spatial and temporal dimensions based on a given query. Datasets like Vid-STG [148], HC-STVG [149], Ego4D-MQ and Ego4D-NLQ [150] are proposed to aid the training and testing for this task.

### v8

@@ I-A1 Conventional Methods
In the early stages of video understanding, handcrafted feature extraction techniques such as Scale-Invariant Feature Transform (SIFT) [lindeberg2012scale], Speeded-Up Robust Features (SURF) [bay2008speeded], and Histogram of Oriented Gradients (HOG) [dalal2005histograms] were used to capture key information in videos. Background Subtraction [sobral2014comprehensive], optical flow methods [tu2022optical], and Improved Dense Trajectories (IDT) [wang2013action, shu2015action] were used to model the motion information for tracking. Since videos can be viewed as time series data, temporal analysis techniques such as Hidden Markov Models (HMM) [liu2003video] have also been used to understand video content. Before the popularity of deep learning, basic machine learning algorithms such as Support Vector Machines (SVM) [sidenbladh2004detecting], Decision Trees [yuan2002automatic], and Random Forests were also used in video classification and recognition tasks. Cluster analysis [chan2008modeling] for classifying video segments, or Principal Component Analysis (PCA) [bouwmans2014robust, hazelhoff2008video] for data dimensionality reduction have also been commonly used methods for video analysis.
@@ I-A2 Early Neural Video Models
Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities. DeepVideo [KarpathyCVPR14] and [earliest] were early methods introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted feature method due to the inadequate use of motion information. Two-stream networks [feichtenhofer2016convolutional] combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted [yue2015beyond]. Temporal Segment Network (TSN) [wang2016temporal] was also designed for long-form video understanding by analyzing and aggregating video segments. Besides TSN, Fisher Vectors (FV) encoding [sekma2015human], Bi-Linear encoding [diba2017deep], and Vector of Locally Aggregated Descriptors (VLAD) [mironicua2016modified] encoding were introduced [li2022transvlad]. These methods improved performance on the UCF-101 [soomro2012ucf101] and HMDB51 [kuehne2011hmdb] datasets. Unlike two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) [tran2015learning]. Inflated 3D ConvNets (I3D) [carreira2017quo] utilizes the initialization and the architecture of 2D CNN, Inception [szegedy2015going], to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) [kay2017kinetics] and Something-Something [goyal2017something] datasets to evaluate the model’s performance in more challenging scenarios. ResNet [he2016deep], ResNeXt [Xie_2017_CVPR], and SENet [Hu_2018_CVPR] were also adapted from 2D to 3D, resulting in the emergence of R3D [Hara_2017_ICCV], MFNet [Chen_2018_ECCV], and STC [Diba_2018_ECCV]. To improve the efficiency, the 3D convolution has been decomposed into cascade 2D and 1D convolution in various studies (e.g., S3D [S3D], ECO [zolfaghari2018eco], P3D [qiu2017learning]). LTC [LTC], T3D [T3D], Non-local [non-local], and V4D [zhang2019v4d] focus on long-form temporal modeling, while CSN [tran2019video], SlowFast [feichtenhofer2019slowfast], and X3D [feichtenhofer2020x3d] tend to attain high efficiency. The introduction of Vision Transformers (ViT) [dosovitskiy2021an] promotes a series of prominent models (e.g., TimeSformer [bertasius2021space], VidTr [li2021vidtr], ViViT [arnab2021vivit], MViT [fan2021multiscale]).
@@ I-A3 Self-supervised Video Pretraining
Transferability [li2023cross, wang2023feature] in self-supervised pretraining models [zhang2023rethinking] for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ requirements for extensive task-specific data. VideoBERT [sun2019videobert] is an early attempt to perform video pretraining. Based on the bidirectional language model BERT [kenton2019bert], pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pretrained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the “pretraining-finetuning” paradigm, many studies on pretrained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT [zhu2020actbert], SpatiotemporalMAE [feichtenhofer2022masked], OmniMAE [girdhar2023omnimae], VideoMAE [tong2022videomae], MotionMAE [yang2022self]) or training strategies (MaskFeat [wei2022masked], VLM [xu2021vlm], ALPRO [li2022align], All-in-One transformer [moritz2020all], MaskViT [gupta2022maskvit], CLIP-ViP [xue2022clip], Singularity [lei2022revealing], LF-VILA [sun2022long], EMCL [jin2022expectation], HiTeA [ye2023hitea], CHAMPAGNE [han2023champagne]).
@@ I-A4 Large Language Models for Video Understanding
Recently, large language models (LLMs) have advanced rapidly [lyu2023gpt]. The emergence of large language models pretrained on extensive datasets has introduced a novel in-context learning capability [zhang2023dnagpt]. This allows them to handle various tasks using prompts without the need for fine-tuning. ChatGPT [OpenAIChatGPT] is the first groundbreaking application built on this foundation. This includes capabilities like generating code and invoking tools or APIs of other models for their use. Many studies are exploring using LLMs like ChatGPT to call vision models APIs to solve the problems in the computer vision field, including Visual-ChatGPT [wu2023visual]. The advent of instruct-tuning has further enhanced these models’ ability to respond effectively to user requests and perform specific tasks [liu2023visual]. LLMs integrated with video understanding capabilities offer the advantage of more sophisticated multimodal understanding, enabling them to process and interpret complex interactions between visual and textual data. Similar to their impact in Natural Language Processing (NLP) [zhao2023survey], these models act as more general-purpose task solvers, adept at handling a broader range of tasks by leveraging their extensive knowledge base and contextual understanding acquired from vast amounts of multimodal data. This allows them to not only understand visual content but also reason about it in a way that is more aligned with human-like understanding. Many works also explore using LLMs in video understanding tasks, namely, Vid-LLMs.
@@ I-B Related Surveys
Previous survey papers either study specific sub-tasks in the area of video understanding or focus on methodologies beyond video understanding. For example, [li2023multimodal] surveys multimodal foundation models for general vision-language tasks, which includes both image and video applications. [survey_vid_cap] and [survey_vid_act_reg] focus on surveying video captioning and action recognition tasks, respectively. Other video understanding tasks, such as the video question answering and grounding, are not considered. Moreover, [survey_vdm], [annepaka2024large], and [zhao2023survey] survey video-related methodologies, such as video diffusion models and LLMs, lacking the concentration on video understanding. [madan2024foundation] centers primarily on video foundation models, with insufficient attention given to language model-based approaches. Despite the significant value to the community, previous survey papers leave a gap in surveying the general video understanding task based on large language models. This paper fills this gap by comprehensively surveying the video understanding task using large language models.
@@ II-A1 Abstract Understanding Tasks
  • Video Classification & Action Recognition: Video classification and action recognition classify videos based on class labels or activities and events categories within a video sequence. Datasets specifically introduced for these tasks include UCF-101 [soomro2012ucf101], HMDB51 [kuehne2011hmdb], Hollywood [hollywood2], ActivityNet [caba2015activitynet], Charades [sigurdsson2016hollywood], Kinetics-400 [kay2017kinetics], Kinetics-600 [carreira2018short], Kinetics-700 [carreira2019short], SomethingSomethingV2 [goyal2017something], HACS [zhao2019hacs], YouTube8M [abuelhaija2016youtube8m], and PortraitMode-400 [han2024video_portraitmode-400]. Usually, Top-K accuracy is adopted as the main metric for these tasks.
  • Text-Video Retrieval: Text-video retrieval task matches and retrieves relevant video clips based on the similarity between video clips and the input textual descriptions. Datasets like Kinetic-GEB [wang2022geb+], MSRVTT [xu2017video], DiDeMo [anne2017localizing], YouCook2 [zhou2018towards], and Youku-mPLUG [xu2023youku] are relevant to this task. The standard evaluation metric for this task is Recall at K (R@K), which measures the accuracy of the first K retrieved results.
  • Video-to-Text Summarization: Video-to-text summarization is a task that generates concise textual summaries of videos. The video summarization approaches are trained to extract and interpret key visual and audio content to produce coherent and informative summaries. ViTT [huang2020multimodal], VideoXum [lin2023videoxum], VideoInstruct-100K [maaz2023video], and Instrcut-V2Xum [hua2024v2xum] are datasets that related to the task. Metrics of BLEU, METEOR, CIDEr, and ROUGE-L often evaluate this task.
  • Video Captioning: Video captioning generates descriptive and coherent textual captions of given videos. The video caption models usually use visual and auditory information from video to produce accurate and contextually relevant descriptions. Notable datasets for this tasks are MSVD [chen2011collecting], MSR-VTT [xu2016msr], TGIF [li2016tgif], Charades [sigurdsson2016hollywood], Charades-Ego [sigurdsson2018charades], YouCook2 [zhou2018towards], Youku-mPLUG [xu2023youku], VAST-27M [chen2023vast], and VideoInstruct-100K [maaz2023video]. This task is often evaluated by metrics of BLEU, METEOR, CIDEr, and ROUGE-L.
  • Video QA: Video Question-Answering (VQA) aims to answer textual questions based on a given video, where the model analyzes visual and auditory information, understands the context, and eventually generates accurate responses. Datasets involved in the QA task are VCR [zellers2019recognition], MSVD-QA [xu2017video], MSRVTT-QA [xu2017video], TGIF-QA [jang2017tgif], Pororo-QA [kim2017deepstory], TVQA [lei2018tvqa], ActivityNet-QA [yu2019activitynet], and NExT-QA [NExT-QA]. This task is evaluated using Top-1, Top-K accuracy.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content. F1-score, Spearman, and Kendall usually evaluate this task as metrics. Commonly-used datasets include SumMe [gygli2014creating], TVSum [song2015tvsum], Ads-1k [tang2022multi], VideoXum [lin2023videoxum], Instrcut-V2Xum [hua2024v2xum].
@@
  • Video Highlight Detection: Video highlight detection aims at identifying and extracting the most important and interesting segments from a video. Commonly used datasets on this task include the YouTube Highlights [youtubeHighlights], the TV-Sum [song2015tvsum], and the Videos Titles in the Wild (VTW) [VTWdataset].
@@
  • Temporal Action/Event Localization: This task aims at identifying the precise temporal segments of actions or events within a video. By analyzing sequential frames, models trained for this task must indicate when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [THUMOS14], ActivityNet-1.3 [caba2015activitynet], and UnAV-100 [geng2023dense_unav100].
@@
  • Temporal Action Proposal Generation: Temporal action proposal generation involves generating candidate segments within a video that are likely to contain actions or events. Relevant datasets like THUMOS’14 [THUMOS14], ActivityNet [caba2015activitynet], and Charades [sigurdsson2016hollywood] are used for the training and evaluation for this task.
@@
  • Video Temporal Grounding: Video temporal grounding is the task of locating specific moments or intervals within a video that correspond to a given textual query. This process involves aligning linguistic descriptions with visual content, enabling precise identification of relevant segments for applications in video search and content analysis. Common benchmarks are Charades-STA [Gao_2017_ICCV], ViTT [huang2020multimodal], DiDeMo [anne2017localizing], and PU-VALOR [tang2024avicuna]. The metrics of R1@0.5 and R1@0.7 often evaluate this task.
@@
  • Moment Retrieval: Moment retrieval is the task of identifying and extracting precise video segments that correspond to a given textual or visual query, which aligns semantic content between queries and video frames. DiDeMo [anne2017localizing] is a dataset for this task.
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying certain frames in a video where significant changes occur and splitting videos based on different events or activities. Kinetics-GEBD [shou2021generic] is a widely-used dataset for this task.
@@
  • Generic Event Boundary Captioning & Grounding: Generic event boundary captioning and grounding involve identifying and describing the transition points between significant events in a video, where Kinetics-GEB+ [wang2022geb+] is the dataset for this task.
@@
  • Dense Video Captioning: Dense video captioning [shao2022region, Wang_2021_ICCV, shao2023textual, long2023capdet, shao2024dcmstrd] aims at generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [krishna2017dense], VidChapters-7M [yang2023vidchapters], YouCook2 [zhou2018towards], and ViTT [huang2020multimodal].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking aims at continuously identifying and following the position of specific objects within a video over time. A good tracking model should maintain accurate and consistent trajectories of objects, even for videos with occlusions, changes in appearance, and motions. Benchmarks like OTB [OTBbenchmark], UAV [UAVbenchmark], and VOT [VOTbenchmark] are commonly used for this task.
  • Re-Identification: Re-Identification (ReID) is the task of recognizing and matching individuals or objects across different video frames or camera views. Common datasets in ReID are Market-1501 [Market1501], CUHK03 [cuhk03], MSMT17 [wei2018person], and DukeMTMC-reID [dukeMTMCreID].
  • Video Saliency Detection: Video saliency detection aims at identifying the most visually important and attention-grabbing regions in a video [moskalenko2024aim]. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [wang2018revisiting_dhf1k], Hollywood-[sigurdsson2016hollywood], UCF-Sports [ucfsport], AVAD [avad_sal], Coutrot1 [coutrot1_2014saliency], Coutrot2 [coutrot2_2016multimodal], ETMD [ETMD_koutras2015perceptually], and SumMe [gygli2014creating].
  • Video Object Segmentation: Video object segmentation aims at partitioning a video into segments that correspond to individual objects, accurately delineating their boundaries over time. YouTube-VOS [youtubevos] and DAVIS [perazzi2016benchmark] are datasets related to this task.
  • Video Instance Segmentation: Video instance segmentation is the task of identifying, segmenting, and tracking each unique object instance within a video. YouTube-VIS [youtubevis] and Cityscapes-Seq [cityscape] are two common benchmarks for this task.
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [ding2023mevis] is a common benchmark for this task.
  • Spatiotemporal Grounding: Spatiotemporal grounding aims to identify and localize specific objects or events within a video’s spatial and temporal dimensions based on a given query. Datasets like Vid-STG [zhang2020does], HC-STVG [hc-stvg-tang2021human], Ego4D-MQ and Ego4D-NLQ [grauman2022ego4d] are proposed to aid the training and testing for this task.